A phanton of 498nm was emitted from a silicon atom. Calculate the energy of all the Atomic levels of silicon and show that the atomic level is quantized


1
Expert's answer
2021-08-23T03:08:06-0400

In our case, mp=0.56m0 (m0=9.1×10-31 kg), T=300 K


Using (3) we got:


gT (E) = 4.08×1023 m-3 × 498




Answer:


2.011×1024 m-3